OpenStudy (anonymous):

ADD OR SUUBTRACT.SHOW WORK!!! 7 2/5 - 4 1/3

OpenStudy (anonymous):

7 2/5=(7*5+2)/5=37/5

OpenStudy (anonymous):

Similarly what is 4 1/3

OpenStudy (anonymous):

HERO HELP ME!!!! I KNOW U KNOW HOW TO DO THIS!!!

hero (hero):

(7 - 4) + (2/5 - 1/3) (7 - 4) + (6/15 - 5/15) 3 + 1/15 3 + 1/15

OpenStudy (anonymous):

whats the anwser?

hero (hero):

lol

OpenStudy (anonymous):

7 2/5 -4 1/3 37/5-13/3 {37*3-13*5}/15 46/15 3 1/15

OpenStudy (anonymous):

But I would prefer to go with @Hero method